# Scale-dependent chirality as a smoking gun for Abelian gauge fields during inflation

### v1

@@ 1 Introduction
A period of accelerated expansion in the very early universe can be readily implemented by means of a scalar field equipped with a sufficiently flat potential [1]. The most constraining data to date on the physics of inflation comes from observations at the largest scales, those of the CMB. The inflationary mechanism has been spectacularly successful in explaining how quantum fluctuations, magnified by the expansion, gave rise to the observed CMB temperature anisotropies and planted the seeds for those matter inhomogeneities that evolved into the cosmos we observe today. Remarkably, all the information gathered from cosmological probes still places only rather qualitative bounds on the microphysics of inflation. Key questions such as those on the nature of the inflationary particle content remain unanswered.
The unprecedented array of cosmological probes set to become operational in the next two decades holds the potential to address such fundamental questions. Probes of the cosmic microwave background such as CMB-S3 and later LiteBIRD and CMB-S4 will improve by almost two order of magnitudes existing bounds on primordial gravitational waves (PGW). Astronomical surveys (such as Euclid, LSST) will close in on the key observable for inflationary (self) interactions, the non-linear parameter $f_{\rm NL}$, and in particular hit the target value $\sigma_{f_{\rm NL}}\sim 1$, a crucial threshold vis-à-vis the single vs multi-field inflation dichotomy. Probes of the 21cm background promise an $\sigma_{f_{\rm NL}}$ further improved by up two orders of magnitude [2] down the line.
Perhaps the most exciting opportunities will be afforded by current and upcoming gravitational wave probes at intermediate and small scales. These have opened up an entire new window (spanning multiple decades in frequency!) on the physics of the early universe. Since LIGO has become operational one may no longer assume a blank canvas for the PGW signal at small frequencies. Further important input is due to pulsar timing arrays. In this context one ought mention the possibility that the recently detected signal [3, 4, 5, 6] in the nHz range may indeed be of primordial origin [7, 8, 9, 4, 10, 11, 12, 13]. In the space of a few years the canvas will be populated by the sensitivity curves of several pulsar timing array experiments and GW interferometers. Undoubtedly then, this is the ideal time to put our most compelling inflationary models to the test, focusing in particular on their the PGW spectrum.
@@
As was recently shown in [28], in inflationary models that contain CS type axion-Abelian gauge field interactions, GW chirality can exhibit non-trivial frequency dependence. This is not due to the standard transition from the vacuum-dominated part of the signal to the range where the sourced contribution takes the lead, but rather to a more subtle non-linear effect in the 1-loop sourcing of the GW power spectrum. The blue-tilted part of a peak of the GW spectrum corresponds to a rapidly increasing gauge quanta production whose sourcing of primordial gravitational waves is nearly insensitive (see section 4.2) to the specific polarization. This is because the sourcing is dominated by large momenta gauge modes which do not discriminate the GW polarization they produce. For a nearly constant or decreasing rate of gauge quanta production the GW spectrum will instead receive the main contribution from the domain of the 1-loop phase-space integral that is most dependent on the polarization $\lambda$ and will therefore exhibit the largest chirality.
The main purpose of this work is to explore this subtle effect for the case of spectator axions coupled to Abelian gauge fields and clarify how this effect is present already in the small back-reaction regime. Furthermore, we show how it can have important implications for the predicted chirality within reach of current and next generation GW detectors. We also compute the power spectrum of scalar perturbations [^2] In our computation we disregard the scalar perturbations of the metric, as typically done in the literature. This is justified in light of e.g. [29] which suggests that the metric perturbations only affect the sourced power spectrum for far smaller values of the particle production parameter than those we consider here. fully numerically, exploiting the numerical scheme of [28]. We use the numerical solutions for both tensor and scalar spectra as benchmarks to evaluate the validity of the semi-analytical WKB approximation commonly employed in analysing these scenarios (see also [30] for a recent study on cross-correlations).
@@
This paper is organized as follows: in section 2 we briefly sumarize the model under study. We give a detailed account of the WKB approximation which is the usual method employed to produce semi-analytic results for the power spectra and explain how the standard procedure can yield inaccurate results for the UV and IR tails of the total power spectra. We end the section by suggesting an alternative and improved approach, still within the WKB approximation scheme. In section 3 we tackle the power spectra using three levels of approximation. We use the fully numerical method as a benchmark and confront these results with the ones arising from the standard WKB procedure as well as our improved method introduced in section 2. Next, in section 4 we explore a few example scenarios whose signatures could be within reach of upcoming experiments. In this context, we establish a relationship between the tilt of the GW spectrum and its expected chirality. Finally, we show that our results and approximations are valid already in the small back-reaction regime. We offer our conclusions in 5. We present some technical details on the normalization of the modes in the improved WKB scheme in appendix A and expand on the fully numerical computation in appendix B.
@@ 3.1 The power spectra of cosmological perturbations
However, as anticipated in section 2.2, this approach leads to inaccurate results since the true profile (see the black curve in Fig. 2) of gauge field fluctuations is considerably different from the Gaussian approximation presented in (2.12). In Fig. 4 we underscore this fact by comparing the scale dependence of $f_{2,j}$ in both approaches using a representative choice of the model parameters. It is clear that, although the Gaussian form (3.8) fits the numerical result in the WKB approach very well in the vicinity of the peaks, it fails to accurately capture especially the IR behavior. Another important lesson we learn from Fig. 4 is the suppression of the net chirality of tensor modes in the IR, where the solid curves in the left panels overlap. This is a consequence of the fact that the phase space for the production of soft tensor modes from an enhanced net $\lambda=-$ vector is dominated by large wavelength modes which are insensitive to the chirality of the gravitational waves produced, as emphasized in [28] (See e.g. their Appendix E).
The results derived in this section have potentially important phenomenological implications. First of all, if the rolling of the axion occurs sufficiently early during inflation, it can generate an observable chiral GW background that dominates over the vacuum at CMB scales. This opens up the possibility of testing even a relatively low scale realization of inflation [47, 48]. However, the amplitude of the sourced GW signal is constrained by the well-established bounds on the scalar perturbations: these can be quite large given that they are sourced by the very same mechanism that generates tensor perturbations (see e.g. [47, 62]). It becomes then crucial to fully characterize sourced tensor perturbations at large scales focusing on observables sensitive to the chirality (e.g. ${\rm TB,EB,BBB}$) of the signal [65, 66], as well as the scale dependence and amplitude of scalar perturbations (e.g. ${\rm TT,TTT,TBB,EBB,BTT,BEE}$) [47, 57].
@@ 4 Phenomenology at intermediate and small scales
In what follows we shall assume that the spectator axion reaches its maximal velocity (and, correspondingly, its leading effect on the observables) far after the time when CMB scales exit the horizon. It follows that the dominant contribution to the scalar and tensor spectra at large scales will be due to vacuum fluctuations. Compliance with CMB observations [68] then places constraints on the inflaton potential $V_{\phi}(\phi)$. In this work we shall be agnostic about the explicit form of the potential whilst ensuring that the slow-roll parameters of the model lead to a scalar power spectrum and spectral tilt that are fully compatible with CMB measurements.
@@ 4.1 Enhanced curvature power spectrum and PBHs
Another interesting phenomenological consequence PBH forming inflationary models is the inevitable production of a stochastic gravitational wave (GW) background [74, 75]. The enhanced spectrum of curvature fluctuations required to produce PBHs acts also as a GW source [76]. The characterization of such GW background (see e.g. [77, 78, 79, 80]) along with the properties of its scalar sources (see e.g. [81, 55]) is an important avenue for testing inflationary models such the ones we study in the present work. We do not pursue this path here as typically the scalar-induced GW signal is overshadowed by the GW production due to vector fields [55]. We refer the interested reader to [82] for a recent comprehensive review on scalar-induced GW backgrounds. In the next section, we focus on the GW background due to the excited gauge quanta of section 3.
@@ 4.2 Chiral GW background from vector field sources
Figure 6: Total gravitational wave spectrum (solid black) together with its right (red dashed) and left (orange dashed) polarized sourced components generated by axion-gauge field dynamics. The dotted curves corresponds to the GW spectrum calculated via the WKB approximated gauge mode functions. Power law integrated sensitivity curves [83] of various current and forthcoming GW probes are shown by shaded regions. The dashed gray line represents the vacuum GW background in the absence of particle production events.
@@
with $\Omega_{r,0}h^{2}\simeq 4.2\times 10^{-5}$ is the radiation density today. In Fig. 6 we present the GW background (4.4) in the frequency $f=k/2\pi$ domain relevant for current and forthcoming GW probes for the two scenarios we have been studying. In these plots, solid and dashed curves refers to the fully numeric approach whereas dotted lines correspond to the GW signal obtained by adopting the WKB approximated gauge mode functions.
@@
The chirality of GWs around the peak. The scale-dependence of the chirality of the GW signal is a very distinct, unique, signature for the class of inflationary models we have been studying. To emphasize this point, we study the frequency dependence of the chirality for the two scenarios analysed in Fig.  6. The quantity we are interested in is the chirality parameter:
@@
with we provide $c_{p}$ coefficients in Table 3. Notice that the sign and absolute value of the fitting coefficients are very close to each other implying a universal behaviour shared by two models for a chirality range $0.3\lesssim\Delta\chi\lesssim 1$. The attentive reader may wonder about the universality of Eq. (4.7) for different choices of the parameter $\delta$, which essentially parametrizes the width of the GW signal (and the mass of the axion around its global minimum) in the vicinity of the peak. We expect this relation to hold, up to order one changes in the $c_{p}$ coefficients, for different values of $\delta$ so long as we remain in the $\delta\ll 3$ regime, where we are nevertheless confined in light of the requirement that the axion be slowly rolling. Naturally, a lower bound on $\delta$ emerges from requiring $\xi$ not be constant (this corresponds to $m_{\chi}\lesssim H$), which would mean a fully chiral signal as per the original analysis of [89]. We expect our consistency relation to be especially useful when used within these two qualitative thresholds.
@@
The validity of the small back-reaction regime. Our derivations for the scale dependent templates of the total power spectra (and their properties) assume that the back-reaction of vector field fluctuations on the background evolution of the spectator axion is negligible. On the other hand, as can be seen from Fig. 2, the gauge field amplification needs to be sufficiently strong to leave observable effects at small scales (see Figs. 5 and 6). It is therefore necessary to have a well-defined criterion to justify our small back-reaction assumption. Identifying such criterion is important also in light of the highly non-trivial dynamics that has been described in the strong back-reaction regime of axion inflation [90, 91, 92, 93, 42, 94, 95, 73, 28, 96, 97]. Furthermore, lattice simulations [98] have shown that in the strong back-reaction regime the fields spatial gradients can become large and have a dramatic impact on the background fields evolution [^11] This is certainly the case at least for a certain range of values of the coupling constant $\lambda$..
The above results above have largely been derived for the case of an axion-inflation. However, it is reasonable to assume that analogous conclusions apply for the spectator case we study in the present work. The mode-by-mode approach we employed disregards the spatial dependence of the axion, and therefore we limit our analysis in the “safe” small back-reaction regime for which the spatial gradient of the axion do not play an important role. In such a regime, the mode-by-mode treatment captures the full dynamics of the system.
@@
We numerically compute $\mathcal{B}$ for parameter choices we adopt in Figs. 5-8. For both models we consider, the resulting normalized spectral back-reaction for various moments in time is shown in Fig. 9. The upper limit (in $x_{*}=k/k_{*}$ space) of the normalized spectral back-reaction curves in the figures are given by the threshold between stable and unstable modes $k_{\rm thr}\equiv-\lambda a\dot{\phi}/(Hf)\Rightarrow x^{\rm thr}_{*}=2\xi(\tau)/(\tau/\tau_{*})$. The threshold is increasing over time (as $\tau/\tau_{*}\to 0$), hence the different upper limits to the back-reaction integral in $\ln x_{*}-{\rm space}$ at different times. We have plotted the normalized spectral back-reaction for 13 different moments in time in 1 e-folding increments symmetrically distributed around $\tau=\tau_{*}$. One notable feature is that there is a delay effect between the moment of largest particle production parameter $\xi=\xi_{*}$ and greatest back-reaction. Additionally, the maximally enhanced mode is shifted (towards the UV) with respect to the mode ($x_{*}=k/k_{*}=1$) that exits the horizon exactly when effective coupling parameter peaks, $\xi(\tau_{*})=\xi_{*}$. This is another manifestation of the delay effect studied in depth in [42].

### v2

@@ 1 Introduction
A period of accelerated expansion in the very early universe can be readily implemented by means of a scalar field equipped with a sufficiently flat potential [1]. The most constraining data to date on the physics of inflation comes from observations at the largest scales, those of the CMB. The inflationary mechanism has been spectacularly successful in explaining how quantum fluctuations, magnified by the expansion, gave rise to the observed CMB temperature anisotropies and planted the seeds for those matter inhomogeneities that evolved into the cosmos we observe today. Remarkably, all the information gathered from cosmological probes still places only rather qualitative bounds on the micro-physics of inflation. Key questions such as those on the nature of the inflationary particle content remain unanswered.
The unprecedented array of cosmological probes set to become operational in the next two decades holds the potential to address such fundamental questions. Probes of the cosmic microwave background such as CMB-S3 [2] and later LiteBIRD [3] and CMB-S4 [4, 5] will improve by almost two order of magnitudes existing bounds on primordial gravitational waves (PGW). Astronomical surveys (such as Euclid [6], LSST [7]) will close in on the key observable for inflationary (self) interactions, the non-linear parameter $f_{\rm NL}$ [8], and in particular hit the target value $\sigma_{f_{\rm NL}}\sim 1$, a crucial threshold vis-à-vis the single vs multi-field inflation dichotomy. Probes of the 21cm background promise an $\sigma_{f_{\rm NL}}$ further improved by up two orders of magnitude [9] down the line.
Perhaps the most exciting opportunities will be afforded by current and upcoming gravitational wave probes at intermediate [10, 11, 12] and small [13, 14, 15, 16, 17, 18] scales. These have opened up an entire new window (spanning multiple decades in frequency!) on the physics of the early universe. Since LIGO has become operational one may no longer assume a blank canvas for the PGW signal at small frequencies. Further important input is due to pulsar timing arrays. In this context one ought mention the possibility that the recently detected signal [11, 19, 20, 21] in the nHz range may indeed be of primordial origin [22, 23, 24, 19, 10, 25, 26, 27]. In the space of a few years the canvas will be populated by the sensitivity curves of several pulsar timing array experiments and GW interferometers. Undoubtedly then, this is the ideal time to put our most compelling inflationary models to the test, focusing in particular on their the PGW spectrum.
@@
As was recently shown in [42], in inflationary models that contain CS type axion-Abelian gauge field interactions, GW chirality can exhibit non-trivial frequency dependence. This is not due to the standard transition from the vacuum-dominated part of the signal to the range where the sourced contribution takes the lead, but rather to a more subtle non-linear effect in the 1-loop sourcing of the GW power spectrum. The blue-tilted part of a peak of the GW spectrum corresponds to a rapidly increasing gauge quanta production whose sourcing of primordial GWs is nearly insensitive (see section 4.2) to the specific polarization. This is because the sourcing in this case is dominated by gauge modes exhibiting relatively large momenta for which the phase space available for the inverse decay process $\delta A+\delta A\to h_{\lambda}$ is identical for both GW polarization states $\lambda=\pm$. For a nearly constant or decreasing rate of gauge quanta production, the GW spectrum will instead receive the main contribution from the domain of the 1-loop phase-space integral that is most dependent on the polarization and will therefore exhibit the largest chirality.
The main purpose of this work is to explore this subtle effect for the case of spectator axions coupled to Abelian gauge fields and clarify how this effect is present already in the small back-reaction regime [^2] See also [43] for recent work focusing on the directly coupled axion inflaton - Abelian gauge field system.. Furthermore, we show how it can have important implications for the predicted chirality within reach of current and next generation GW detectors. We also compute the power spectrum of scalar perturbations [^3] In our computation we disregard the scalar perturbations of the metric, as typically done in the literature. This is justified in light of e.g. [44] which suggests that the metric perturbations only affect the sourced power spectrum for far smaller values of the particle production parameter than those we consider here. fully numerically, exploiting the numerical scheme of [42]. We use the numerical solutions for both tensor and scalar spectra as benchmarks to evaluate the validity of the semi-analytical WKB approximation commonly employed in analysing these scenarios (see also [45] for a recent study on cross-correlations).
@@
This paper is organized as follows: in section 2 we briefly summarize the model under study. We give a detailed account of the WKB approximation which is the usual method employed to produce semi-analytic results for the power spectra and explain how the standard procedure can yield inaccurate results for the UV and IR tails of the total power spectra. We end the section by suggesting an alternative and improved approach, still within the WKB approximation scheme. In section 3 we tackle the power spectra using three levels of approximation. We use the fully numerical method as a benchmark and confront these results with the ones arising from the standard WKB procedure as well as our improved method introduced in section 2. Next, in section 4 we explore a few example scenarios whose signatures could be within reach of upcoming experiments. In this context, we establish a relationship between the tilt of the GW spectrum and its expected chirality. Finally, we show that our results and approximations are valid already in the small back-reaction regime. We offer our conclusions in 5. We present some technical details on the normalization of the modes in the improved WKB scheme in appendix A and expand on the fully numerical computation in appendix B.
@@ 3.1 The power spectra of cosmological perturbations
However, as anticipated in section 2.2, this approach leads to inaccurate results since the true profile (see the black curve in Fig. 2) of gauge field fluctuations is considerably different from the Gaussian approximation presented in (2.12). In Fig. 4 we underscore this fact by comparing the scale dependence of $f_{2,j}$ in both approaches using a representative choice of the model parameters. It is clear that, although the Gaussian form (3.8) fits the numerical result in the WKB approach very well in the vicinity of the peaks, it fails to accurately capture especially the IR behavior. Another important lesson we learn from Fig. 4 is the suppression of the net chirality of tensor modes in the IR, where the solid curves in the left panels overlap. This is a consequence of the fact that the phase space for the production of soft tensor modes is dominated by the enhanced gauge fields that exhibit relatively large loop momenta. As a result, large wavelength ($x_{*}=k/k_{*}\ll 1$) sourced tensor perturbations are comparable for the two GW polarization (left/right) states. This leads to the suppression of net chirality of the produced GWs in the IR, as first emphasized in [42]. For the sake of completeness and clarity, we have now added a brief technical discussion on these important points in Appendix B. The results derived in this section have potentially important phenomenological implications. First of all, if the rolling of the axion occurs sufficiently early during inflation, it can generate an observable chiral GW background that dominates over the vacuum at CMB scales. This opens up the possibility of testing even a relatively low scale realization of inflation [62, 63]. However, the amplitude of the sourced GW signal is constrained by the well-established bounds on the scalar perturbations: these can be quite large given that they are sourced by the very same mechanism that generates tensor perturbations (see e.g. [62, 77]). It becomes then crucial to fully characterize sourced tensor perturbations at large scales focusing on observables sensitive to the chirality (e.g. ${\rm TB,EB,BBB}$) of the signal [80, 81], as well as the scale dependence and amplitude of scalar perturbations (e.g. ${\rm TT,TTT,TBB,EBB,BTT,BEE}$) [62, 72].
@@ 4 Phenomenology at intermediate and small scales
In what follows we shall assume that the spectator axion reaches its maximal velocity (and, correspondingly, its leading effect on the observables) far after the time when CMB scales exit the horizon. It follows that the dominant contribution to the scalar and tensor spectra at large scales will be due to vacuum fluctuations. Compliance with CMB observations [83] then places constraints on the inflaton potential $V_{\phi}(\phi)$ [^10] See e.g. [84] for a recent work in the context of natural inflation that exhibit direct CS type coupling with gauge fields.. In this work we shall be agnostic about the explicit form of the potential whilst ensuring that the slow-roll parameters of the model lead to a scalar power spectrum and spectral tilt that are fully compatible with CMB measurements.
@@ 4.1 Enhanced curvature power spectrum and PBHs
Another interesting phenomenological consequence PBH forming inflationary models is the inevitable production of a stochastic GW background [90, 91]. The enhanced spectrum of curvature fluctuations required to produce PBHs acts also as a GW source [92]. The characterization of such GW background (see e.g. [93, 94, 95, 96]) along with the properties of its scalar sources (see e.g. [97, 70]) is an important avenue for testing inflationary models such the ones we study in the present work. We do not pursue this path here as typically the scalar-induced GW signal is overshadowed by the GW production due to vector fields [70]. We refer the interested reader to [98] for a recent comprehensive review on scalar-induced GW backgrounds. In the next section, we focus on the GW background due to the excited gauge quanta of section 3.
@@ 4.2 Chiral GW background from vector field sources
Figure 6: Total GW spectrum (solid black) together with its right (red dashed) and left (orange dashed) polarized sourced components generated by axion-gauge field dynamics. The dotted curves corresponds to the GW spectrum calculated via the WKB approximated gauge mode functions. Power law integrated sensitivity curves [99] of various current and forthcoming GW probes are shown by shaded regions. The dashed gray line represents the vacuum GW background in the absence of particle production events.
@@
with $\Omega_{r,0}h^{2}\simeq 4.2\times 10^{-5}$ is the radiation density today [^12] Note that our expression assumes standard radiation domination for the post inflationary universe and we are neglecting changes to the effective number of degrees of freedom as they are irrelevant for $T\geq 100\;{\rm GeV}$ which corresponds to $f>4\cdot 10^{-4}\;{\rm Hz}$ and only mildly affect PTA scale GWs by a factor 2 enhancement.. In Fig. 6 we present the GW background (4.4) in the frequency $f=k/2\pi$ domain relevant for current and forthcoming GW probes for the two scenarios we have been studying. In these plots, solid and dashed curves refers to the fully numeric approach whereas dotted lines correspond to the GW signal obtained by adopting the WKB approximated gauge mode functions.
@@
The chirality of GWs around the peak. The scale-dependence of the chirality of the GW signal is a very distinct, unique signature of the class of inflationary models we have been studying. To emphasize this point, we study the frequency dependence of the chirality for the two scenarios analysed in Fig.  6. The quantity we are interested in is the chirality parameter:
@@
Figure 9: Evolution of scale GW signal from IR to UV scales (using log-frequency scale) in the $1-\Delta\chi$ vs $n_{\rm gw}$ plane. Consistency condition (4.7) is shown by solid black curve, parametrizing the evolution of the spectral tilt $n_{\rm gw}$ in terms of the excess chirality as the GW signal rise and falls around its peak value (shown by the red cross) where $n_{\rm gw}=0$. Overlapping bands are shown to indicate the frequency range where the signal is above the sensitivity curves of forthcoming GW probes we consider in Figure 6.
@@
with we provide $c_{p}$ coefficients in Table 3. Notice that the sign and absolute value of the fitting coefficients are very close to each other implying a universal behaviour shared by two models for a chirality range $0.3\lesssim\Delta\chi\lesssim 1$. We illustrate these points in Fig. 9 where the evolution of the GW signal in the $1-\Delta\chi$ vs $n_{\rm gw}$ plane is shown for both scenarios we consider in Fig. 6. We observe that far in the IR, chirality is vanishing ($\Delta\chi\sim 0$ or $1-\Delta\chi\sim 1$) while $n_{\rm gw}\lesssim 0$. As the vector field sources become active, chirality $\Delta\chi$ of the GW signal grows along the positive $n_{\rm gw}$ branch while maintaining its growth for frequencies following the peak where $n_{\rm gw}<0$. Eventually particle production shuts down and hence the GW signal returns to the origin of the “phase space” where $\Delta\chi\to 0$ ($1-\Delta\chi\to 1$) and $n_{\rm gw}\lesssim 0$. The accuracy of the consistency relation (4.7) around the maximum of the GW signal ($n_{\rm gw}=0$) – shown by the solid dark curve – is visible. Furthermore, the potential coverage of the signal with multiple probes indicated by the colored thick bands together with its parametrization by the consistency relation we derived in this work is particularly encouraging.
The attentive reader may wonder about the universality of Eq. (4.7) for different choices of the parameter $\delta$, which essentially parametrizes the width of the GW signal (and the mass of the axion around its global minimum) in the vicinity of the peak. We expect this relation to hold, up to order one changes in the $c_{p}$ coefficients, for different values of $\delta$ so long as we remain in the $\delta\ll 3$ regime, where we are nevertheless confined in light of the requirement that the axion be slowly rolling. Naturally, a lower bound on $\delta$ emerges from requiring $\xi$ not be constant (this corresponds to $m_{\chi}\lesssim H$), which would mean a fully chiral signal as per the original analysis of [105]. We expect our consistency relation to be especially useful when used within these two qualitative thresholds.
@@
The validity of the small back-reaction regime. Our derivations for the scale dependent templates of the total power spectra (and their properties) assume that the back-reaction of vector field fluctuations on the background evolution of the spectator axion is negligible. On the other hand, as can be seen from Fig. 2, the gauge field amplification needs to be sufficiently strong to leave observable effects at small scales (see Figs. 5 and 6). It is therefore necessary to have a well-defined criterion to justify our small back-reaction assumption. Identifying such criterion is important also in light of the highly non-trivial dynamics that has been described in the strong back-reaction regime of axion inflation [106, 107, 108, 109, 57, 110, 111, 112, 89, 42, 113, 114]. Furthermore, lattice simulations [115] have shown that in the strong back-reaction regime the fields spatial gradients can become large and have a dramatic impact on the background fields evolution [^14] This is certainly the case at least for a certain range of values of the coupling constant $\lambda$..
The results above have largely been derived for the case of an axion-inflation. However, it is reasonable to assume that analogous conclusions apply for the spectator case we study in the present work. The mode-by-mode approach we employed disregards the spatial dependence of the axion, and therefore we limit our analysis in the “safe” small back-reaction regime for which the spatial gradient of the axion do not play an important role. In such a regime, the mode-by-mode treatment captures the full dynamics of the system.
@@
We numerically compute $\mathcal{B}$ for parameter choices we adopt in Figs. 5-8. For both models we consider, the resulting normalized spectral back-reaction for various moments in time is shown in Fig. 10. The upper limit (in $x_{*}=k/k_{*}$ space) of the normalized spectral back-reaction curves in the figures are given by the threshold between stable and unstable modes $k_{\rm thr}\equiv-a\lambda\dot{\phi}/f\Rightarrow x^{\rm thr}_{*}=2\xi(\tau)/(\tau/\tau_{*})$. The threshold is increasing over time (as $\tau/\tau_{*}\to 0$), hence the different upper limits to the back-reaction integral in $\ln x_{*}-{\rm space}$ at different times. We have plotted the normalized spectral back-reaction for 13 different moments in time in 1 e-folding increments symmetrically distributed around $\tau=\tau_{*}$. One notable feature is that there is a delay effect between the moment of largest particle production parameter $\xi=\xi_{*}$ and greatest back-reaction. Additionally, the maximally enhanced mode is shifted (towards the UV) with respect to the mode ($x_{*}=k/k_{*}=1$) that exits the horizon exactly when effective coupling parameter peaks, $\xi(\tau_{*})=\xi_{*}$. This is another manifestation of the delay effect studied in depth in [57].